OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (texaschic101):

-5 | x + 1 | = 10 --- divide both sides by 5 | x + 1 | = 2 x + 1 = 2 -(x + 1) = 2 x = 2 - 1 -x - 1 = 2 x = 1 -x = 2 + 1 -x = 3 x = -3 so x = 1 and x = -3

OpenStudy (anonymous):

Thank you!!!